OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 7, 14, 28, 56, ... Converges; 49 Converges; 105 Diverges Converges; 217

OpenStudy (anonymous):

Do you know how to find the common ratio, r?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

r = (2nd #)/(1st #)

OpenStudy (anonymous):

so it would be 98?

OpenStudy (anonymous):

Then after you have r: If |r| > 1, the series diverges If 0<|r|<1, the series converges to \[S_n=\frac{ a_1 }{ 1-r }\]

OpenStudy (anonymous):

r = 14/7 = 2

OpenStudy (anonymous):

ohhh divide not multiply ok i get it, so it diverges?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

you're welcome